Using your data, calculate the moles and mass of acetylsalicylic acid in each tablet (molar mass = 180.16 g/mol).Data Concentration of NaOH 107 M Solutions Mass of Aspirin Tablet #11 Mass of Aspirin Tablet #21 Volume of NaOH mixed with Ta
Answer #1


Molar mass of Acetyl salicylic acid given as 180.16g/mol

That means 180.16 gm of acetylsalicylic acid is = 1 mol

                     1 gm of acetylsalicylic acid is = 1/180.16 mol

Trial 1 mass 0.161 gm of acetylsalicylic acid is = (1/180.16) \times0.161 mol = 8.94 \times10-4 mol

For trial 2 mass 0.282 gm of acetylsalicylic acid is = (1/180.16) \times0.282 mol = 1.57\times10-3 mol